OpenStudy (anonymous):

how do i solve y=x^2-4x-45 to find all x intercepts

OpenStudy (anonymous):

Factor!

OpenStudy (anonymous):

How do you do that? can you help?

OpenStudy (anonymous):

Well you want to make it into (x+a)(x+b) form

OpenStudy (anonymous):

so when you remultiply it it gives you x^2+ax+bx+ab

OpenStudy (anonymous):

what are some values of a and b that could give you ab=-45?

OpenStudy (anonymous):

9 and 5

OpenStudy (anonymous):

right, note that one of them needs to be negative. Now how would you make ax+bx = -4x from your 9 and 5?

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

is the answer possibly 5 and -9?

OpenStudy (anonymous):

why yes it is. so what does your y=(x+a)(x+b) become?

OpenStudy (anonymous):

(x-5) and (x+9)

OpenStudy (anonymous):

you have your a and b values backwards actually. when you multiply (x-5)(x+9) = x^2+4x-45

OpenStudy (anonymous):

so its -9 and +5?

OpenStudy (anonymous):

yes those are your a and b values. Next you need to find values of x to provide 0=(x-9)(x+5) those will be when your y=0

OpenStudy (anonymous):

NO WAIT -5 and 9